# Version of Tuesday $24^{\text{th}}$ June, 2025

### v1

@@ 4.2. A posteriori verification of minimum: Riccati equation
To verify if the Riccati equation (22) has a bounded solution over $[0,1]$, we propose the following algorithm. Since the linear system is Hamiltonian, we adopt a symplectic integrator, and since the blocks of the matrix $M$ are available at the grid points found while solving the BVP, we adopt a scheme that only requires this information. Below, we refer to $\widehat{\mathbf{x}}(t_{k})$ as the computed trajectory at the gridpoints.
@@
    Compute $A,B,C$ and $\dot{B}$, along the computed trajectory $\widehat{\mathbf{x}}(t_{k})$, at gridpoints $t_{k}$, $k=0,1,\dots,N$, where $t_{0}=0$, $t_{N}=1$. To compute $\dot{B}$, we use forward differences: $\dot{B}(t_{k})=\frac{B(t_{k+1})-B(t_{k})}{t_{k+1}-t_{k}}$, so that we have $\dot{B}$ at all gridpoints except the last one. Then, form the blocks $P=C$ and $Q=A-\frac{\dot{B}+\dot{B}^{T}}{2}$ at the grid points.
@@
    Approximate the solution of (22) by using the symplectic Euler scheme on the linear system identified in Lemma 3.10. With $U_{0}=I$ and $V_{0}=0$, compute
    $\begin{cases}V_{k+1}&=V_{k}-h_{k}Q(t_{k})U_{k}\\ U_{k+1}&=U_{k}-h_{k}P^{-1}(t_{k})V_{k+1}\end{cases}\quad h_{k}=t_{k+1}-t_{k},\,\ k=0,1,\dots,N-1\ .$
    NB: obviously, no inverse is explicitly formed. Namely, we do not explicitly form $P^{-1}$, but merely solve a linear system; that is, we find
    $U_{k+1}=U_{k}-h_{k}Z_{k}\ ,\,\ \text{where}\,\ P_{k}Z_{k}=V_{k+1}\ .$
    NNB: We do not need to form the approximation to (22) $W_{k+1}=V_{k+1}U_{k+1}^{-1}\$, since all we care about is whether or not there are conjugate points, but not the solution of (22).
  • (iii)
    In step (ii), we monitor if $\det(U_{k+1})$ becomes negative, in order to rule out having missed singularities. At the beginning, $\det(U_{0})=1$.
  • (iv)
    If there is a change of sign in the determinant of $U_{k+1}$ (step (iii) above), then we declare that there are conjugate points in $[0,1]$, otherwise we do not. If there are no conjugate points, then the extremal we found is a minimizer.
4.3. Discrete OT
@@
    In the linear assignment problem, given the cost matrix $C$, we match a row to a column so to minimize the cost. To illustrate, in (2), take $k_{0}=k_{1}=k$ and $\mu_{i}=\frac{1}{k}$, $\nu_{j}=\frac{1}{k}$, $\forall i,j$. In this case, $\pi$ will need to be a non-negative doubly stochastic matrix up to a scalar multiplication. Popular solution techniques for the assignment problem are the Hungarian algorithm (see [12, 14, 8, 16]), relaxation algorithms like the auction algorithm of Bertsekas and variations of it ([2, 17]), as well as Dijkstra algorithm, [6], which is what we used in our numerical experiments in Section 5 as implemented in the builtin function matchpairs of Matlab, see [7]. We note that –with this technique– we cannot split masses.
@@
    Sinkhorn. This relatively recent technique, [5], has won the favor of many people, because of the flexibility to handle the general setup of (2). First, one regularizes the minimization problem by adding to it an entropic regularization term to smooth out the problem (actually, one adds the negative of entropy regularization). That is, for a fixed $\epsilon>0$, one seeks the minimum of $J_{\epsilon}$ below:
@@
    An appealing feature of the method is that it is simple to implement an algorithm to find the optimal plan for the regularized problem (25). Defining ${\mathbf{u}}\in\mathbb{R}^{k_{0}}$, ${\mathbf{v}}\in\mathbb{R}^{k_{1}}$, and $K\in\mathbb{R}^{k_{0}\times k_{1}}$ with $K_{ij}=e^{-\frac{c({\mathbf{x}}_{i},{\mathbf{y}}_{j})}{\epsilon}}$, to recover the optimal plan $\pi^{*}_{\epsilon}$ of (25), for fixed $\epsilon>0$, one needs to perform a sufficient number of iterations on the vectors $\mathbf{u}$} and $\mathbf{v}$} to converge to $\pi^{*}_{\epsilon}$ (see [5]):
@@
In principle, as $\epsilon$ decreases, the minimum value of $J_{\epsilon}$ in (25) approaches the minimum value of $J$ in (3), but the convergence is slow: convergence rate is linear and the contraction constant is $\mathcal{O}(1-e^{-c/\epsilon})$. Moreover, in practice, there is always some value $\epsilon_{0}$, which depends on the problem, such that even for $0<\epsilon$, but $\epsilon<\epsilon_{0}$, the iterations in (26) will not converge due to numerical instabilities created by the $\frac{1}{\epsilon}$ term. So, some trade-off is necessary between the choice of $\epsilon$ and the accuracy with which we approximate the solution of (3); see Section 5 for values of $\epsilon$ we used in practice.
@@ 5.1. Optimal Trajectory
Here we show computation of optimal trajectories for the energy and length minimization problems in the context of two-point configurations. Our scope here is to verify that we satisfy both necessary and sufficient conditions for the energy-minimizing solution and also to show that the curves obtained by energy-minimization or length-minimization are indeed the same curve. In the numerical examples below, we used $21$ equispaced homotopy steps (see Section 4.1). In all of our numerical experiments below, we used absolute and relative error tolerances of $10^{-4}$.
The following numerical examples in 2-d and 3-d are examined:
@@
(E2) ${\mathbf{a}}=\left(\begin{smallmatrix}-7\\ -7\end{smallmatrix}\right)\ ,\ {\mathbf{b}}=\left(\begin{smallmatrix}7\\ 7\end{smallmatrix}\right)\ ,\ K({\mathbf{x}})=\sin(x_{1}(t))-\sin(x_{2}(t))+3\ ,$
@@
(a) Energy solution, cost is $2.2917$
(b) Length solution, cost is $2.1409$
@@
$\left|L({\mathbf{z}})^{2}-2E({\mathbf{y}})\right|=1.95\times 10^{-4},{\qquad\text{and}\qquad}\left|L({\mathbf{z}})-L({\mathbf{y}})\right|=1.84\times 10^{-6}.$
(a) Energy solution, cost is $1410.8$
(b) Length solution, cost is $53.119$
@@
(a) $\det(U)$ values, $\min_{k}\det(U_{k})=1$
(b) $K({\mathbf{y}})\|\dot{\mathbf{y}}\|$ values, range is $2.04*10^{-7}$
@@
$\left|L({\mathbf{z}})^{2}-2E({\mathbf{y}})\right|=6.27\times 10^{-8}{\qquad\text{and}\qquad}\left|L({\mathbf{z}})-L({\mathbf{y}})\right|=1.58\times 10^{-8}.$
(a) Energy solution, cost is $1.9684$
(b) Length solution, cost is $1.9841$
@@ 5.2. Discrete OT Examples
First, we consider three problems, (E4, E5, E6), formulated on a uniform mesh grid consisting of 9 points in the two-dimensional case and 8 points in the three-dimensional case with uniform masses. The masses are assumed to be uniform, with $\mu_{i}=\frac{1}{k}$, $\nu_{j}=\frac{1}{k}$, $\forall i,j$, where $k$ represents the number of points. To compute each entry of the cost matrix, 5 equispaced homotopy steps were used. In rare cases where the boundary value problem solver did not attain the specified tolerance, the number of homotopy steps was increased to a maximum of 51. For the sake of completeness, we point out that each entry of the cost matrix is computed by solving the relevant Euler-Lagrange equation, (12) or (14); for the cost given by (5), verification of optimality of all trajectories needed to setup the cost matrix was always successful.
@@
(E5) $\begin{split}&\mathbf{\Omega_{X}}=[-5,-4]\times[-4,-3]\ ,\ \mathbf{\Omega_{Y}}=[3,4]\times\bigg[\frac{11}{4},\frac{15}{4}\bigg]\ ,\\ &K({\mathbf{x}})=\sin(x_{1}(t))-\sin(x_{2}(t))+3\ ,\end{split}$
@@ 5.2.1. Optimal Assignment Algorithm
In Figure 7, we see that the assignment plans obtained for the energy-based and length-based cost matrices in Example (E4) are the same. However, the assignment plans for Examples (E5) and (E6) differ, as shown in Figures 89.
(a) Energy solution, total cost is $2.8792$
(b) Length solution, total cost is $2.3982$
@@
(a) Energy solution, total cost is $439.17$
(b) Length solution, total cost is $29.615$
Figure 8. Optimal assignment for Example (E5)
(a) Energy solution, total cost is $2.0153$
(b) Length solution, total cost is $2.0052$
@@ 5.2.2. Linear Programming Formulation: Sinkhorn Method
When using this formulation, beside the aforementioned difference in the transport solutions when using length or energy costs, we also witness a possible mass splitting, as expected. Figures 1012 illustrate the optimal transport plans for Examples (E4)–(E6). In these visualizations, the brightness of each line corresponds to the amount of mass transported along the respective path. As shown in Figures 10(a)12(a), the transport plans derived from the energy-based cost matrix are sparser and more closely resemble an assignment mapping. In contrast, when using the length-based cost matrix, the Sinkhorn algorithm produces denser transport matrix. Consequently, the transported mass is more diffusely distributed among the target points, as evident in Figures 10(b)12(b). These observations further underscore the nonlinear relationship between the cost matrices, whose entries are computed via either energy or length minimization.
@@
(a) Energy solution, total cost is $2.8809$
(b) Length solution, total cost is $2.4$
@@
(a) Energy solution, total cost is $439.47$
(b) Length solution, total cost is $29.625$
Figure 11. Sinkhorn solution for Example (E5), $\epsilon=\frac{3}{4}$
(a) Energy solution, total cost is $2.0156$
(b) Length solution, total cost is $2.0068$
@@ 6. Conclusions
In this work, our main effort has been directed toward forming the cost matrix underpinning the solution of a discrete optimal transport problem in a nonuniform environment. From the mathematical modeling point of view, the non-uniformity of the environment reflects in having that the cost of moving one unit of mass from location $\mathbf{a}$} to location $\mathbf{b}$}, call it $c({\mathbf{a}},{\mathbf{b}})$, is a weighted version of the standard length or energy costs; see (4) and (5). As a consequence, there is no close expression for $c({\mathbf{a}},{\mathbf{b}})$, and $c({\mathbf{a}},{\mathbf{b}})$ must be found as the minimizer among all possible paths joining $\mathbf{a}$} and $\mathbf{b}$}. For this latter task, we considered, and solved, the associated Euler-Lagrange equations. We further provided verifiable sufficient conditions of optimality of the path that we find when solving the Euler-Lagrange equations. We proposed new algorithms, and gave evidence of their effectiveness on several numerical examples in 2 and 3 space dimensions.

### v2

@@ 3.2. A posteriori verification of minimum: Riccati equation
To verify if the Riccati equation (23) has a bounded solution over $[0,1]$, we propose the following simple algorithm. Since the blocks of the matrix $M$ are available at the grid points found while solving the BVP, we adopt a scheme that only requires this information. Below, we refer to $\widehat{\mathbf{x}}(t_{k})$ as the computed trajectory at the gridpoints.
@@
    Let $\widehat{\mathbf{x}}(t_{k})$ be the computed trajectory at gridpoints $t_{k}$, $k=0,1,\dots,N$, where $t_{0}=0$, $t_{N}=1$. Form $A,B,C$ and $\dot{B}$, at $t_{k}$, $k=0,1,\dots,N-1$. To compute $\dot{B}$, we use forward differences: $\dot{B}(t_{k})=\frac{B(t_{k+1})-B(t_{k})}{t_{k+1}-t_{k}}$, so that we have $\dot{B}$ at all grid points except the last one. Then, we form the blocks $P$, $Q=A-\frac{\dot{B}+\dot{B}^{T}}{2}$ and $E=\frac{B-B^{T}}{2}$ at all grid points, except $t_{N}$.
@@
    Approximate the solution of (23) by using Euler method on the linear system identified in Theorem 2.10. That is, with $U_{0}=I$ and $V_{0}=0$, compute
    $\begin{split}&\begin{bmatrix}U_{k+1}\\ V_{k+1}\end{bmatrix}\ =\ \bigl(I+h_{k}M(t_{k})\bigr)\begin{bmatrix}U_{k}\\ V_{k}\end{bmatrix}\ ,\,\,\text{where}\\ &h_{k}=t_{k+1}-t_{k},\,\ k=0,1,\dots,N-1,\,\ t_{0}=0\ ,\,t_{N}=1.\end{split}$
    Of course, no inverse $P^{-1}$ is explicitly computed when forming $M$, and a linear system is instead solved. To clarify, we have
    $\begin{split}&U_{k+1}\ =\ U_{k}+h_{k}\left[-(P^{-1}E^{T})(t_{k})U_{k}+P^{-1}(t_{k})V_{k}\right]\,\ \text{and}\\ &V_{k+1}\ =\ V_{k}+h_{k}\left[(Q-EP^{-1}E^{T})(t_{k})U_{k}+EP^{-1}(t_{k})V_{k}\right]\ ,\end{split}$
    and so we solve the two systems $PX=E^{T}$ and $PY=V_{k}$.
  • (ii)
    Observe that we do not need to form the approximation to (23) $W_{k+1}=V_{k+1}U_{k+1}^{-1}\$, since all we care about is whether or not there are conjugate points, but not the solution of (23). So, we monitor if $\det(U_{k+1})$ becomes negative, in order to rule out having missed singularities. At the beginning, $\det(U_{0})=1$. If there is a change of sign in the determinant of $U_{k+1}$ (step (ii) above), then we declare that there are conjugate points in $[0,1]$, otherwise we do not. If there are no conjugate points, then the extremal we found is a minimizer.
3.3. Solving the discrete OT
@@
    In the linear assignment problem, given the cost matrix $C$, we match a row to a column so to minimize the cost. To illustrate, in (3), take $k_{0}=k_{1}=k$ and $\mu_{i}=\frac{1}{k}$, $\nu_{j}=\frac{1}{k}$, $\forall i,j$. In this case, $\pi$ will need to be a non-negative doubly stochastic matrix up to a scalar multiplication. Popular solution techniques for the assignment problem are the Hungarian algorithm (see [kuhn1955hungarian, munkres1957algorithms, edmonds1972theoretical, tomizawa1971some]), relaxation algorithms like the auction algorithm of Bertsekas and variations of it ([bertsekas1981new, walsh2019real]), as well as Dijkstra algorithm, [Dijkstra], which is what we used in our numerical experiments in Section 4 as implemented in the builtin function matchpairs of Matlab, see [DuffKoster]. We note that –with this technique– we cannot split masses.
@@
    Sinkhorn. This relatively recent technique, [Cuturi2013a], has won the favor of many people, because of the flexibility to handle the general setup of (3). First, one regularizes the minimization problem by adding to it an entropic regularization term to smooth out the problem (actually, one adds the negative of entropy regularization). That is, for a fixed $\epsilon>0$, one seeks the minimum of $J_{\epsilon}$ below:
@@
    An appealing feature of the method is that it is simple to implement an algorithm to find the optimal plan for the regularized problem (25). Defining ${\mathbf{u}}\in\mathbb{R}^{k_{0}}$, ${\mathbf{v}}\in\mathbb{R}^{k_{1}}$, and $K\in\mathbb{R}^{k_{0}\times k_{1}}$ with $K_{ij}=e^{-\frac{c({\mathbf{x}}_{i},{\mathbf{y}}_{j})}{\epsilon}}$, to recover the optimal plan $\pi^{*}_{\epsilon}$ of (25), for fixed $\epsilon>0$, one needs to perform a sufficient number of iterations on the vectors $\mathbf{u}$} and $\mathbf{v}$} to converge to $\pi^{*}_{\epsilon}$ (see [Cuturi2013a]):
@@
In principle, as $\epsilon$ decreases, the minimum value of $J_{\epsilon}$ in (25) approaches the minimum value of $J$ in (4) at a linear rate. However, in practice, there is always some value $\epsilon_{0}$, which depends on the problem, such that for $\epsilon<\epsilon_{0}$, there are numerical stability issues; because of the $\frac{1}{\epsilon}$ term; see [PeyreCuturi] and [Schmitzer] for a discussion of this aspect, and see Section 4 for the values of $\epsilon$ we used in our examples.
@@ 4.1. Optimal Trajectory
Below, we show computation of optimal trajectories for the energy and length minimization problems in the context of two-point configurations. Our scope here is to verify that we satisfy both necessary and sufficient conditions for the energy-minimizing solution and also to show that the curves obtained by energy-minimization or length-minimization are indeed the same curve. In the numerical examples below, we used $21$ equispaced homotopy steps (see Section 3.1) and in all of our experiments we used absolute and relative error tolerances of $10^{-4}$ for solving the BVP. Table 1 reports the optimal transport cost for each problem discussed below, together with the total computation time, on a fixed mesh of $10^{4}$ time-steps. From Table 2, we note that the computation time of each homotopy step is roughly the same, which confirms that the numerical method is effectively working with a fixed mesh.
The following test cases in 2-d and 3-d are considered:
@@
(E2) ${\mathbf{a}}=\left(\begin{smallmatrix}-7\\ -5\end{smallmatrix}\right)\ ,\ {\mathbf{b}}=\left(\begin{smallmatrix}6\\ 7\end{smallmatrix}\right)\ ,\ K({\mathbf{x}})=\sin(x_{1}(t))-\sin(x_{2}(t))+3\ ,$
@@
Energy Cost Length Cost
Example Solution (6) Time Solution (5) Time
(E1) $2.2917$ $23.63$ sec. $2.1409$ $20.422$ sec.
(E2) $1108.4$ $23.566$ sec. $47.082$ $23.345$ sec.
(E3) $1.9684$ $24.796$ sec. $1.9841$ $23.458$ sec.
Table 1. Minimal cost values and computation time
Energy Cost Length Cost
Homotopy Step (E1) (E2) (E3) (E1) (E2) (E3)
$\alpha=0$ $0.3659$ $0.45082$ $0.55204$ $0.42017$ $0.50373$ $0.41298$
$\alpha=0.1$ $1.1119$ $1.4195$ $1.1878$ $0.97861$ $1.0408$ $1.131$
$\alpha=0.2$ $1.1061$ $1.2395$ $1.2002$ $1.0344$ $1.2453$ $1.0631$
$\alpha=0.3$ $1.1176$ $1.2487$ $1.2254$ $1.0008$ $1.2670$ $1.1905$
$\alpha=0.4$ $1.1128$ $1.2448$ $1.213$ $0.95345$ $1.2678$ $1.1444$
$\alpha=0.5$ $1.1201$ $1.2477$ $1.234$ $0.95799$ $1.2678$ $1.1459$
$\alpha=0.6$ $1.2048$ $1.0189$ $1.2004$ $0.95029$ $1.0441$ $1.1423$
$\alpha=0.7$ $1.2048$ $1.0241$ $1.2017$ $1.02$ $1.0424$ $1.1704$
$\alpha=0.8$ $1.1836$ $1.0355$ $1.1817$ $1.0197$ $1.0451$ $1.1794$
$\alpha=0.9$ $1.1707$ $1.0319$ $1.1827$ $1.0337$ $1.0549$ $1.1902$
$\alpha=1$ $1.3944$ $1.0280$ $1.2024$ $1.0323$ $1.0484$ $1.2687$
Total Time: $\mathbf{23.63}$ $\mathbf{23.566}$ $\mathbf{24.796}$ $\mathbf{20.422}$ $\mathbf{23.345}$ $\mathbf{23.458}$
Table 2. Computation time (in seconds) across 21 homotopy steps
@@
(a) Energy solution
(b) Length solution
@@
$\left|L({\mathbf{z}})^{2}-2E({\mathbf{y}})\right|=1.24\times 10^{-5}\ ,{\qquad\text{and}\qquad}\left|L({\mathbf{z}})-L({\mathbf{y}})\right|=1.32\times 10^{-7}\ .$
(a) Energy solution
(b) Length solution
@@
(a) $\det(U)$ values, $\min_{k}\det(U_{k})=0.99995$
(b) $K({\mathbf{y}})\|\dot{\mathbf{y}}\|$ values, range is $3.87*10^{-7}$
@@
$\left|L({\mathbf{z}})^{2}-2E({\mathbf{y}})\right|=6.27\times 10^{-8}\ ,{\qquad\text{and}\qquad}\left|L({\mathbf{z}})-L({\mathbf{y}})\right|=1.58\times 10^{-8}\ .$
(a) Energy solution
(b) Length solution
@@ 4.2. Discrete OT Examples
Consider two problems, (E4 and E5), formulated on a uniform mesh grid consisting of 9 points in the two-dimensional case and 8 points in the three-dimensional case with uniform masses. The masses are assumed to be uniform, with $\mu_{i}=\frac{1}{k}$, $\nu_{j}=\frac{1}{k}$, $\forall i,j$, where $k$ represents the number of points. To compute each entry of the cost matrix, 5 equispaced homotopy steps were used. In rare cases where the boundary value problem solver did not attain the specified tolerance, the number of homotopy steps was increased to a maximum of 51. For the sake of completeness, we point out that each entry of the cost matrix is computed by solving the relevant Euler-Lagrange equation, (12) or (14) and that for the cost given by (6), verification of optimality of all trajectories needed to setup the cost matrix was always successful.
@@ 4.2.1. Optimal Assignment Algorithm
In Figure 7, we see that the assignment plans obtained for the energy-based and length-based cost matrices in Example (E4) are the same. However, the assignment plan for Example (E5) differ, as shown in Figure 8.
(a) Energy solution, total cost is $2.9975$
(b) Length solution, total cost is $2.4478$
@@
(a) Energy solution, total cost is $2.0574$
(b) Length solution, total cost is $2.0262$
@@ 4.2.2. Linear Programming Formulation: Sinkhorn Method
When using this formulation, beside the aforementioned difference in the transport solutions when using length or energy costs, we also witness a possible mass splitting, as expected. Figure 9 and Figure 10 illustrate the optimal transport plans for Examples (E4) and (E5). In these visualizations, the brightness of each line corresponds to the amount of mass transported along the respective path. As shown in Figures 9(a) and 10(a), the transport plans derived from the energy-based cost matrix are sparser and more closely resemble an assignment mapping. In contrast, when using the length-based cost matrix, the Sinkhorn algorithm produces denser transport matrix. Consequently, the transported mass is more diffusely distributed among the target points, as evident in Figures 9(b) and 10(b). These observations further underscore the nonlinear relationship between the cost matrices, whose entries are computed via either energy or length minimization.
@@
(a) Energy solution, total cost is $2.9996$
(b) Length solution, total cost is $2.4490$
@@
(a) Energy solution, total cost is $2.0577$
(b) Length solution, total cost is $2.0278$
@@ 5. Conclusions
In this work, our main effort has been directed toward forming the cost matrix underpinning the solution of a discrete optimal transport problem in a nonuniform environment. From the mathematical modeling point of view, the non-uniformity of the environment reflects in having that the cost of moving one unit of mass from location $\mathbf{a}$} to location $\mathbf{b}$}, call it $c({\mathbf{a}},{\mathbf{b}})$, is a weighted version of the standard length or energy costs; see (5) and (6). As a consequence, there is no close expression for $c({\mathbf{a}},{\mathbf{b}})$, and $c({\mathbf{a}},{\mathbf{b}})$ must be found as the minimizer among all possible paths joining $\mathbf{a}$} and $\mathbf{b}$}. For this latter task, we considered, and solved, the associated Euler-Lagrange equations. We further provided verifiable sufficient conditions of optimality of the path that we find when solving the Euler-Lagrange equations. We proposed and implemented our algorithms, and gave evidence of their effectiveness and reliability on several numerical examples in 2 and 3 space dimensions.
Declarations
– Ethical approval: Not applicable.
– Availability of supporting data: Not applicable.
– Funding: Not applicable.
– Competing interests: The authors declare no competing interests.
– Authors’ contributions: L.D. and D.O. contributed equally to the manuscript..
– Acknowledgements: Not applicable.